### v1

@@ 6 Conclusion
In this paper, we introduced RelGNN, a novel graph neural network framework specifically designed to address the structural inefficiencies of existing heterogeneous GNNs for relational databases. By leveraging atomic routes—which capture high-order tripartite structures—we designed a composite message passing mechanism that enables direct single-hop interactions between heterogeneous nodes. This avoids redundant aggregation and mitigates information entanglement, leading to more efficient and accurate predictive modeling. Through extensive evaluation on RelBench, a diverse benchmark covering 30 predictive tasks across seven relational databases, RelGNN consistently outperforms state-of-the-art baselines, achieving up to a 25% improvement in predictive accuracy. Our findings emphasize the limitations of conventional heterogeneous GNNs when applied to relational data and highlight the necessity of models that explicitly account for primary-foreign key relationships.
Impact Statement
@@ Appendix B Visualization of Primary-Foreign Key Relationships
In this section, we visualize the primary-foreign key relationships of all datasets in RelBench (rel-f1 is visualized in Figure  1).
Figure 4: The primary-foreign key relationship of rel-amazon dataset.
Figure 5: The primary-foreign key relationship of rel-hm dataset.
Figure 6: The primary-foreign key relationship of rel-avito dataset.
Figure 7: The primary-foreign key relationship of rel-event dataset.
Figure 8: The primary-foreign key relationship of rel-stack dataset.
Figure 9: The primary-foreign key relationship of rel-trial dataset.

### v2

@@ 7 Conclusion
In this paper, we introduced RelGNN, a novel graph neural network framework specifically designed to address the structural inefficiencies of existing heterogeneous GNNs for relational databases. By leveraging atomic routes, we designed a composite message passing mechanism that enables direct single-hop interactions between the source and destination nodes. This avoids redundant aggregation and highlights key predictive signals, leading to more efficient and accurate predictive modeling. Through extensive evaluation on RelBench, a diverse benchmark covering 30 predictive tasks across 7 relational databases, RelGNN outperforms state-of-the-art baselines on the vast majority of tasks, achieving up to a 25% improvement in predictive accuracy. Our findings emphasize the limitations of conventional heterogeneous GNNs when applied to relational data and highlight the necessity of models that explicitly account for primary-foreign key relationships.
@@ Appendix B Visualization of Primary-Foreign Key Relationships
We visualize the primary-foreign key relationships of all datasets in RelBench in Figure 4 (rel-f1 is visualized in Figure 1 (a)).
(a) rel-amazon
(b) rel-hm
(c) rel-avito
(d) rel-event
(e) rel-stack
(f) rel-trial
Figure 4: Primary-foreign key relationships of datasets in RelBench.
Appendix C Ablation Study
To isolate the contribution of atomic routes, we conduct an ablation study comparing variants of RelGNN with and without the attention mechanism. In our instantiation and main experiments, the FUSE operation of RelGNN is instantiated using the same design as GraphSAGE (cf. Equation (4)). For this ablation, we additionally configure the AGGR operation to use GraphSAGE-style aggregation instead of attention. This ensures that the only architectural difference between the baseline and RelGNN is the inclusion of atomic routes.
As shown in Tables CC, and C, RelGNN achieves consistent gains over the baselines, even when using the same aggregation mechanism. Notably, the performance improvement is not attributable to attention alone: the gap between RelGNN with and without attention is modest. These results indicate that the performance gains stem primarily from the use of atomic routes, which enable more efficient and targeted message passing across relational structures.
We also include additional baselines using GAT (Veličković et al., 2018) and GIN (Xu et al., 2019) as backbone architectures. As shown in Tables CC, and C, RelGNNoutperforms all baselines regardless of the backbone GNN.
Table 5: Ablation results for entity classification (ROC-AUC(%), higher is better) on RelBench test set. The best values are in bold, and the second-best values are underlined.
[…]
Table 6: Ablation results for entity regression results (MAE, lower is better) on RelBench test set. The best values are in bold, and the second-best values are underlined.
[…]
Table 7: Ablation results for recommendation results (MAP(%), higher is better) on RelBench test set. The best values are in bold, and the second-best values are underlined.
[…]
Appendix D Additional Discussion
Scalability and Efficiency in Large Relational Databases.
A common concern is whether the complexity of computing atomic routes increases with database size. In our framework, atomic routes are computed at the schema graph level, where nodes represent table types rather than individual entities. Even in large databases with millions of rows, the number of table types typically remains in the tens. This schema-level design makes RelGNN agnostic to the number of entities and allows it to scale efficiently as the database grows.
Another related concern is the complexity introduced by tables with many foreign keys. In practice, we follow the standard RDL sampling strategy, which samples a fixed-size local neighborhood around each seed node. As a result, only a subset of the foreign key links is activated at training or inference time. The number of active connections remains bounded by the sampling configuration, ensuring that the message passing remains efficient even when some tables connect to many others.
A Potential Approach for Handling Self-Loop Tables.
Although self-loop tables are rare in relational databases (only 1 out of 7 datasets in RelBench), it is worthwhile to improve the model handling of such cases. The key to addressing this challenge is enhancing the model’s ability to distinguish messages coming from the same table type (for self-loop tables) versus those from different types (for non-self-loop tables). To address this, we propose incorporating relative positional encoding (RPE) over the schema graph (where nodes are table types and edges are primary-foreign key relations). RPE helps the model identify whether two tables are of the same type and also offers additional benefits. Since the schema graph defines the database structure at a macro level, it can provide global relational context that complements the local structural information captured by GNNs on the relational entity graph. Moreover, schema graphs are typically small (with only tens of nodes even in large databases) and static, so the RPE can be precomputed once with negligible overhead and reused throughout training and inference.
To evaluate, we implemented an RPE method based on  Huang et al. : eigen-decompose the Laplacian of the schema graph $L=V\text{diag}(\lambda)V^{\top}$, apply $m$ element-wise MLPs $\phi_{k}(\cdot)$ to obtain $Q[:,:,k]=V\text{diag}(\phi_{k}(\lambda))V^{\top}$, for $k\in[m]$, resulting in a tensor $Q\in\mathbb{R}^{N\times N\times m}$, and project via an MLP $\rho$ to obtain $RPE=\rho(Q)\in\mathbb{R}^{N\times N\times d}$, where $N$ is the number of node-types of $d$ is the dimension of node embeddings. For a message from node-type $i$ to $j$, we update the original message $m$ as $m^{\prime}=m+\alpha\cdot RPE[i,j]$ with learnable $\alpha$. This resulted in a 2% improvement on both classification tasks in rel-stack, which contains self-loops.
Integration with Advanced Components.
We focus on improving the core design of message passing, but the RelGNN framework is flexible and accommodates more advanced components.
  • Advanced GNN Aggregators. RelGNN allows easy integration of different GNN aggregators by re-instantiating the AGGR operation in Equation (4.3). For example, PNA  (Corso et al., 2020) improves expressiveness by combining multiple aggregators and degree-scalers. It can be integrated into RelGNN by instatiating Equation (4.3) with $AGGR(\mathbf{h}_{dst}^{(l)},\{\{\mathbf{h}_{fuse}^{(l)}\}\})=\mathbf{W}_{\text{proj}}\mathbf{h}_{dst}^{(l)}+\sum_{fuse\in\mathcal{N}(dst)}M(\mathbf{h}_{fuse}^{(l)})$ where $M(\cdot)$ is the PNA operator. Model-agnostic aggregators can also be integrated on top of RelGNN. For instance, Jumping Knowledge Network  (Xu et al., 2018) improves performance by adaptively combining representations from multiple layers. This can be implemented by collecting outputs from each RelGNN layer and passing them to PyG’s JumpingKnowledge module, with the result used as input to the final prediction layer.
[…]